AUGER RECOMBINATION DYNAMICS IN ISOLATED MERCURY CLUSTER ANIONS Hg_{n}^{−} (n = 9 − 22) FOLLOWING S → P INTERBAND EXCITATION AT 4.65 eV

\underline{O. T. EHRLER}, G. B. GRIFFIN, R. M. YOUNG, A. KAMMRATH, D. M. NEUMARK, Department of Chemistry, University of California at Berkeley, Berkeley, CA 94720, United States; O. CHESH-NOVSKY, School of Chemistry, The Raymond and Beverly Sackler Faculty of Exact Sciences, Tel Aviv University, Tel Aviv 69978, Israel.

The electronic structure of mercury clusters follows the Bloch-Wilson model over a wide size range. As the number of atoms increases, the s/p-hybridisation is enhanced, accompanied by a widening of the filled valence s band and the empty p-band. This leads to a reduction of the gap between those bands, which will eventually merge. The band-gap can be readily observed by photoelectron spectroscopy of mercury cluster anions Hg_{n}^{−}, and its closure was extrapolated to occur at n ≈ 400 atoms.^{a} At sufficiently high photon energies, absorption competes the photodetachment promoting a second electron into the p-band and leaving a hole of s-character. The excitation can interact both with nuclear and electronic degrees of freedom before recombining and emitting the remaining electron from the p-band. This leads to characterstic tails in the photoelectron spectra as a fingerprint of the preceeding relaxation dynamics.^{b}

We have directly measured the dynamics of electronic relaxation following s to p interband excitation of mass selected mercury cluster anions Hg_{n}^{−} (with n = 9 − 22) using ultrafast time-resolved pump/probe photoelectron spectroscopy.^{c} Auger decay of the excited clusters was found to occur on a timescale of 300 − 500 fs, changing abruptly between n = 12 and 13. These dynamics also define an upper limit of the non-adiabatic coupling and are an order of magnitude faster than results previously reported on such electronic relaxation in Hg_{n}. This difference is interpreted as the result of correlated electron dynamics, and mechanisms are posited for relaxation of both the excited electrons in the p-band and the hole in the s-band.

aB. v. Issendorff, and O. Cheshnovsky, Annu. Rev. Phys. Chem. 56, 549 (2005).

bR. Busani, R. Giniger, T. Hippler, and O. Cheshnovsky, Phys. Rev. Lett. 90, 083401 (2003).

cG. B. Griffin, A. Kammrath, O. T. Ehrler, R. M. Young, O. Cheshnovsky, and D. M. Neumark Chem. Phys. (2007)

doi:10.1016/j.chemphys.2007.12.005.